2 Ca3(PO4)2 + 6 SiO2 + 10C ---> 6 CaSiO3 + 10 CO(g) + P4
How many grams of calcium phosphate would be needed to make 100 grams of phosphorus? How much calcium silicate would be formed as a by-product?
1
Expert's answer
2016-02-03T09:17:26-0500
Question:
2 Ca3(PO4)2 + 6 SiO2 + 10C ---> 6 CaSiO3 + 10 CO(g) + P4
How many grams of calcium phosphate would be needed to make 100 grams of phosphorus? How much calcium silicate would be formed as a by-product?

Solution:
N(P4) = 100 / (30.97 * 4) = 0.8072
M(Ca3(PO4)2) = 0.8072 * 2 * 310.18 = 500.77

Answer: 500.77 g